a car accelerated along straight road from rest to 21 m/s in 5s. what is the car average acceleration?
1
Expert's answer
2020-09-21T08:31:49-0400

The average acceleration tells us how quickly the velocity changed:


"a=\frac{\Delta v}{t}=\frac{v_2-v_1}{t}\frac{21-0}{5}=4.2\text{ m/s}^2."